# Grounding DINO: Marrying DINO with Grounded Pre-Training for Open-Set Object Detection

### v3

@@ Appendix E Comparison between DINO and Grounding DINO
Appendix F Visualizations
Appendix G Marry Grounding DINO with Stable Diffusion
Grounding DINO L finishes the detection task in Fig. 1. The results in Fig. 1 (b) are generated by two processes. First, we detect objects with Grounding DINO. Then we generate masks by masking out the detected objects or backgrounds. After that, we feed original images, image masks, and generation prompts to the stable diffusion models [42] to render new images. We use the releases checkpoints in https://github.com/Stability-AI/stablediffusion for new image generation. More results are available in Figure 8.
Using GLIGEN for Grounded Generation
To enable fine-grained image editing, we combine the Grounding DINO with GLIGEN [GLIGEN]. GLIGEN supports grounding results as inputs and can generate objects on specific positions. We can assign each bounding box an object with GLIGEN, as the cake and dog in Figure 9 (c) (d). Moreover, GLIGEN can full fill each bounding box, which results in better visualization, as that in Figure 9 (a) (b). For example, we use the same generative prompt in Figure 8 (b) and Figure 9 (b). The GLIGEN results ensure each bounding box with an object and fulfills the detected regions.
Figure 8: Combination of Grounding DINO and Stable Diffusion. We first detect objects with Grounding DINO and then perform image inpainting with Stable Diffusion. * The input human face in the row (e) is generated by StyleGAN.
Figure 9: Combination of Grounding DINO and GLIGEN. We first detect objects with Grounding DINO and then perform image inpainting with GLIGEN. * The phrase prompts are separated by semicolons. We assign phrase prompts to bounding boxes randomly.
Appendix H Effects of RefC and COCO Data
We add the RefCOCO/+/g (we note it as “RefC” in tables) and COCO into training in some settings. We explore the influence of these data in Table 13. The results show that RefC helps improve the COCO zero-shot and fine-tuning performance but hurts the LVIS and ODinW results. With COCO introduced, the COCO results is greatly improved. It shows that COCO brings marginal improvements on LVIS and slightly decreases on ODinW.
Model Pre-Train COCO minival LVIS minival ODinW
Zero-Shot Fine-Tune Zero-Shot Zero-Shot
Grounding DINO T O365,GoldG 48.1 57.1 25.6 20.0
Grounding DINO T O365,GoldG,RefC 48.5 57.3 21.9 17.7
Grounding DINO T O365,GoldG,RefC,COCO 56.1 57.5 22.3 17.4
@@ Appendix H Effects of RefC and COCO Data
Appendix I Model Efficiency
We compare the model size and efficiency between Grounding-DINO-T and GLIP-T in Table 14. The results show that our model has a smaller parameter size and better efficiency than GLIP.

### v5

@@ 0.D.1 Detection Visualizations
0.D.2 Physical meaning of language-guided query
We visualize the top 900 queries (boxes) in language-guided query selection across various prompts in Figure 7. It shows our model can have dynamic queries during inference. We will incorporate the analyses into the revised paper.
Figure 7: Top queries in language-guided query selection.
0.D.3 Comparison of RefCOCO and Grounding Data
The RefCOCO has a different formulation with grounded training, resulting in a big performance gap without the RefCOCO dataset. As shown in Fig. 8, each RefCOCO text prompt corresponds to only one box, while our model tends to predict multiple objects.
Figure 8: Our model predictions and ground-truths in RefCOCO.
0.D.4 Marry Grounding DINO with Stable Diffusion for Object Detection and Inpainting
We present an image editing application in Fig. 1 (b). The results in Fig. 1 (b) are generated by two processes. First, we detect objects with Grounding DINO and generate masks by masking out the detected objects or backgrounds. After that, we feed original images, image masks, and generation prompts to an inpainting model (typical Stable Diffusion [41]) to render new images. We use the released checkpoints in https://github.com/Stability-AI/stablediffusion for new image generation. More results are available in Figure 9.
The “detection prompt” is the language input for Grounding DINO, while the “generation prompt” is for the inpainting model.
0.D.5 Marry Grounding DINO with Stable Diffusion for Object Detection and Grounded Generation
To enable fine-grained image editing, we combine the Grounding DINO with GLIGEN [27]. We use the “phrase prompt” in Figure 10 as the input phrases of each box for GLIGEN.
GLIGEN supports grounding results as inputs and can generate objects on specific positions. We can assign each bounding box an object with GLIGEN, as shown in Figure 10 (c) (d). Moreover, GLIGEN can full fill each bounding box, which results in better visualization, as that in Figure 10 (a) (b). For example, we use the same generative prompt in Figure 9 (b) and Figure 10 (b). The GLIGEN results ensure each bounding box with an object and fulfills the detected regions.
Figure 9: Combination of Grounding DINO and Stable Diffusion. We first detect objects with Grounding DINO and then perform image inpainting with Stable Diffusion. “Detection Prompt” and “Generation Prompt” are inputs for Grounding DINO and Stable Diffusion, respectively. *The input human face in the row (e) is generated by StyleGAN.
Figure 10: Combination of Grounding DINO and GLIGEN. We first detect objects with Grounding DINO and then perform image inpainting with GLIGEN. “Detection Prompt” and “Generation Prompt” are inputs for Grounding DINO and Stable Diffusion, respectively. “Phrase Prompt” are language inputs for each bounding box. The phrase prompts are separated by semicolons. *We assign phrase prompts to bounding boxes randomly.